Fermi National Laboratory
 

Accelerator Update for July 30, 2007 - August 1, 2007

 

Monday July 30

The day shift began with Operations monitoring the Tevatron (TeV) store 5588 (89.85E30) and the Recycler stash (157.3E10), with the Antiproton Source (Pbar) stacking (48.8E10), with Linac, Booster, and Main Injector (MI) operating normally, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams, and with MiniBooNE and the Meson MTest experiments taking beam. NuMI will remain off until after the summer shutdown.

At 8:37 AM, Operations reported that a TeV vertical separator at sector B11 had sparked. It caused a loss of luminosity.

At 8:58 AM, a Booster vacuum pump (IP41) tripped off and would not reset. Operators contacted an expert who later reported that the pump had a bad power supply, which he replaced.

At 11:39 AM, two Booster RF stations (BRF1 & 3) tripped off. They both reset, but BRF1 tripped off again. It reset. Operators discovered that the trips were due to a particular time line in the Time Line Generator (TLG) that MI experts were using for a multi batch study. The MI experts altered their timeline.

MI experts began a slipstacking study at 12:56 PM.

At 1:43 PM, a Booster power supply (GMPS) tripped off on a ground fault indication. Experts investigated and had the power supply back on by 2:02 PM.

The Neutron Therapy Facility (NTF) reported at 3:44 PM that they had completed treatment on their last patient for the day.

Operations reported at 3:58 PM that the Recycler had stashed antiprotons three times during the day shift.

At 4:10 PM, a Linac Klystron RF station (KRF3) tripped off. It reset. Also, the P2 transfer line tripped off due to a power supply current drift (I:CPSF1I), which was part of its protection system. It reset.

At 4:40 PM, the P2 transfer line tripped off again due to CPSF1I. Operations contacted an expert, who raised the supply’s limit. Operators reset the transfer line.

MTest finished using the beam for the night at 6:07 PM.

At 7:19 PM, BRF11 tripped off. It reset.

Operators prepared the Meson M03-5 areas for an access at 8:50 PM.

At 10:41 PM, LRF3 tripped off for the first of two times in 20 minutes. It reset.

At 10:43 PM, the Linac low energy buncher tripped off. It reset.

At 11:40 PM, LRF3 tripped off again. Operators reseated the temperature/humidity interlock module and this cleared the problem.

Operations reported at 11:58 PM that the Recycler had stashed three times during the evening shift.

Tuesday July 31

The midnight shift began with Operations monitoring store 5588 (33.76E30) and the stash (332.5E10), with Pbar stacking (30.5E10), and with MiniBooNE taking beam.

At 2:55 AM, the Linac buncher tripped off. It reset.

At 4:27 AM, BRF18 tripped off. It reset.

At 5:32 AM, a Pbar Debuncher quadrupole power supply (D:Q724) tripped off. Operators investigated and reported that there was no control power. Operations called in an expert who found and replaced a blown fuse. He had the power supply back on by 6:10 AM.

Operations reported at 7:48 AM that the Recycler had stashed three times during the midnight shift.

Operations reported that the intensity reduction program in preparation for the summer shutdown had begun at 8 AM.

Operations established beam to MTest at 9:03 AM.

Operations terminated store 5588 at 9:18 AM.

Operations put the TeV into shot setup at 9:35 AM.

At 11:03 AM, KRF1 tripped off. It reset.

At 11:05 AM, KRF1 had gone out of phase. It came back.

At 11:17 AM, BRF7 tripped off. It reset.

Recycler experts accessed the MI-31 pit at 11:49 AM.

Operations established store 5589 at 12:04 PM with an initial luminosity of 264.3E30.

Operations reported that a Recycler expert began work on the new damper system at 12:28 PM.

At 2:10 PM, a Pbar sweeper magnet power supply (M:USWA) tripped off. It reset.

At 2:23 PM, Operations reported that a CDF beam monitor voltage supply (B0LMV1) had a drooping output. The problem was that this supply was tied to part of the CDF detector (SVX) and the drooping could cause a beam abort. Trying to raise the voltage could also cause an abort. Experts investigated.

Pbar experts began a reverse proton study at 3:35 PM.

Operations reported at 3:58 AM that the Recycler had stashed two times during the day shift.

At 4:04 PM, the Linac buncher tripped off. It reset.

At 4:40 PM, LRF3 tripped off and wouldn’t reset. An expert investigated and had the station back up by 6:22 PM.

At 6:24 PM, Operations reported that a Pbar Debuncher RF station (DRF1-8) was off due to a blown fuse. Operators searched for a replacement fuse.

At 6:58 PM, KRF5 tripped off. It reset.

At 7:22 PM, Operations reported that the new Recycler vertical dampers were causing the emittances to grow. Operators turned off the dampers and called in an expert.

MTest finished using the beam for the night at 7:40 PM.

Operations reported at 8 PM that turning off the Recycler vertical dampers allowed the emittances to drop to normal levels.

Operations reported at 8:23 PM that a Recycler expert switched back to the old damper system.

Operations reported at 8:27 PM that the DRF1-8 problem wasn’t due to a bad fuse. After consulting with an expert, operators tried various fixes, but none of them worked. Operators did discover that stacking improved with DRF1-1 and DRF1-8 off.

At 8:51 PM, the MiniBooNE beam permit tripped off due to a power supply (HV860) trip. It reset.

RF experts came in to investigate the DRF1 problems at 9:11 PM. 30 minutes later they reported that corrupted files had been loaded into the ramp card. The loaded the proper curves by hand.

Operations reported at 11:57 PM that the Recycler had stashed two times during the evening shift.

Wednesday August 1

The midnight shift began with Operations monitoring store 5589 (80.13E30) and the stash (128.9E10), with Pbar stacking (39.3E10), and with MiniBooNE taking beam.

At 3:59 AM, a Recycler kicker (R:KPS2A) tripped off. It reset.

Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster and MI studies today. Experts will continue to investigate possible solutions to the B0HVLM problem.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/01/2007    email Fermilab